OpenStudy (idealist10):

Use variation of parameters to find a particular solution of x^2*y"+xy'-y=2x^2+2, given that y1=x and y2=1/x satisfy the complementary equation.

OpenStudy (idealist10):

\[y _{p}=u _{1}x+u _{2}\frac{ 1 }{ x }\] \[u'_{1}x+u'_{2}\frac{ 1 }{ x }=0\] \[u'_{1}-u'_{2}\frac{ 1 }{ x^2 }=2x^2+2\] ------------------------------------ \[2u'_{2}\frac{ 1 }{ x }=-2x^3-2x\] \[u'_{2}=-x^4-x^2\] \[u _{2}=-\frac{ x^5 }{ 5 }-\frac{ x^3 }{ 3}\] \[u'_{1}=x^2+1\] \[u _{1}=\frac{ x^3 }{ 3 }+x\] \[y _{p}=\frac{ x^4 }{ 3 }+x^2-\frac{ x^4 }{ 5 }-\frac{ x^2 }{ 3 }\] But the answer in the book says \[y _{p}=\frac{ 2(x^2-3) }{ 3 }\]. What's wrong?

OpenStudy (idealist10):

@freckles

OpenStudy (loser66):

You forgot divided both sides by x^2, so that your f(t) ( the RHS) is not correct. It should be \(2+\dfrac{2}{x^2}\)

OpenStudy (loser66):

This mistake make your \(u_1'-u_2'\dfrac{1}{x^2}=2x^2+2\) becomes wrong.

hartnn (hartnn):

you need to bring x^2*y"+xy'-y=2x^2+2 in the form \( y'' + p(x) y' + q(x) y = g(x) \) only then you'll able to use the method of parameters this way. So, as suggested by Loser, please divide that entire equation by x^2 first

hartnn (hartnn):

\(u'_{1}-u'_{2}\frac{ 1 }{ x^2 }=2+\frac{2}{x^2} \)